OpenStudy (noahbred):

HELP please medal renting a bike cost $30 for the first day and $10 for each additional day write an equation in simplified form to find the total cost C for a bike rental in terms of the number of days d it was rented.

OpenStudy (noahbred):

please help me solve

OpenStudy (anonymous):

Because of the total cost, we want our equation to equal C. We also need to incorporate +30 into the equation because it is a cost that does not vary according to the amount of days that the bike is being rented. Finally, you need to add 10d because 10 is added per day the bike is being rented. Could you write an equation that uses all of these parts?

OpenStudy (noahbred):

C+30+10d+10 is this the equation

OpenStudy (noahbred):

@OrangeMaster

OpenStudy (anonymous):

No quite, you want your equation to equal C. You need to take: 30 10d =C And make an equation using these parts.

OpenStudy (noahbred):

30+10d=C

OpenStudy (anonymous):

Yup, so for example if the bike was rented for 6 day the equation would be: 30+10(6)=90 90 dollars for 6 days, but yes your equation is correct.

OpenStudy (noahbred):

tyyyyyy

OpenStudy (anonymous):

No problem, do you understand, or so you need me to go into further detail?

OpenStudy (anonymous):

*do you

OpenStudy (noahbred):

more detail

OpenStudy (anonymous):

Okay, the bike is going to cost $30 dollars regardless. It does not vary depending on the amount of days it is rented. However, the $10 dollars does vary depending on the amount of days, and that is why we add the variable 'd' beside the $10. So, the amount of days + the initial cost of the bike should equal the total cost of the bike, or 'C'

OpenStudy (noahbred):

so its c=90

OpenStudy (anonymous):

Well, that was just an example, but yes if the bike was rented for 6 days, d=6 and C=$90

OpenStudy (anonymous):

The general equation though is: 30+10d=C

OpenStudy (noahbred):

ohh so if D= 5 then c would be 80?

OpenStudy (noahbred):

Omg Tanks!!!!!!!!

OpenStudy (anonymous):

Yes, because you would take the amount of days that the bike is being rented and input that value in for the variable 'd', and then you would solve accordingly until you got the final cost of 'C'

OpenStudy (anonymous):

No problem.